I have a rectangular garden with an area of 99 square meters. if the with of the garden is 9 meters. what is the perimeter of the garden?
Answer by checkley77(12844)

AREA=LW
99=L*9
99=9L
L=99/9
L=11 ANS. FOR THE LENGTH.
PERIMETER=2L+2W
P=2*11+2*9
P=22+18
P=40 ANS. FOR THE PERIMETER.